find the slope and the y intercept of the graph of -3x+ 2y=5
Answer by venugopalramana(3286)

find the slope and the y intercept of the graph of -3x+ 2y=5
2Y=3X+5
Y=(3/2)X+(5/2).THIS IS IN STD.FORM ......Y=MX+C...WHERE M=SLOPE AND C=Y INTERCEPT
SLOPE = 3/2
Y INTERCEPT =5/2